OpenStudy (anonymous):

Show an equation and a solution for the problem. It takes John 20 hours to paint a building. It takes Sam 15 hours to paint the same building. How long will it take for them to paint the building if they work together, with Sam starting one hour later than John? Round to the nearest tenth.

OpenStudy (ash2326):

John takes 20 hours to paint a building so in one hour he'll paint 1/20 th of the building similarly Sam paints 1/15 th of the building in an hour together they 'll paint in an hour =1/15 +1/20 of the building i.e 35/300 sam paints in an hour 1/15 of building so building left to paint= 1-1/15=14/15 so time taken for the two to paint he rest= (14/15)/(35/300)= 8 hours